A car is moving at a rate of 50 m/s when suddenly the driver hit the horn that



emits 20000Hz frequency. What will be the perceived frequency of a man



standing at the sidewalk if the speed of the sound is 340 m/s and the car is



moving away from the location of the man?

1
Expert's answer
2021-11-19T10:50:53-0500

We can find the frequency observed by the man from the Doppler Shift formula:


"f_o=f_s(\dfrac{v}{v+v_s}),""f_o=2\times10^4\ Hz\times(\dfrac{340\ \dfrac{m}{s}}{340\ \dfrac{m}{s}+50\ \dfrac{m}{s}})=1.74\times10^4\ Hz."